# custom-line = letter = : , command = dashedline , ccommand = cdashedline , tikz = dashed

### v2

@@ Abstract.
copyright: noneconference: ; ;
@@ 1. Introduction
One potential solution that could partially address these challenges is to adapt the “pre-train, prompt” paradigm from natural language processing (NLP) to the graph domain. In NLP, prompt-based learning has effectively generalized pre-trained language models across diverse tasks. For example, a sentiment classification task like “The WebConf will take place in the scenic city of Singapore in 2024” can be reframed by appending a specific textual prompt “I feel so [MASK]” to the end. It is highly likely that a language model pre-trained on next word prediction will predict “[MASK]” as “excited” instead of “frustrated”, without necessitating extensive fine-tuning. With this methodology, certain downstream tasks can be seamlessly aligned with the pre-training objectives. While few prior work (Sun et al., 2023, 2022; Fang et al., 2022a; Liu et al., 2023a; Tan et al., 2023) has delved into crafting various prompting templates for graphs, their emphasis remains strictly on homogeneous graphs. This narrow focus underscores the last challenge inherent to the heterogeneous graph structures typical of the Web:
@@
To comprehensively address all four aforementioned challenges, we propose HetGPT, a general post-training prompting framework tailored for heterogeneous graphs. Represented by the acronym Heterogeneous Graph Prompt Tuning, HetGPT serves as an auxiliary system for HGNNs that have undergone constrastive pre-training. At the core of HetGPT is a novel graph prompting function that reformulates the downstream node classification task to align closely with the pretext contrastive task. We begin with the the virtual class prompt, which generalizes the intricate characteristics of each class in the embedding space. Then we introduce the heterogeneous feature prompt, which acts as a task-specific augmentation to the input graph. This prompt is injected into the feature space and the prompted node features are then passed through the pre-trained HGNN, with all parameters in a frozen state. Furthermore, a multi-view neighborhood aggregation mechanism, that encapsulates the complexities of the heterogeneous neighborhood structure, is applied to the target node, generating a node token for classification. Finally, Pairwise similarity comparisons are performed between the node token and the class tokens derived from the virtual class prompt via the contrastive learning objectives established during pre-training, which effectively simulates the process of deriving a classification decision. In summary, our main contributions include:
@@ 2. Related Work
Heterogeneous graph pre-training. Pre-training techniques have gained significant attention in heterogeneous graph machine learning, especially under the scenario with limited labeled nodes (Liu et al., 2022; Xie et al., 2022). Heterogeneous graphs, with their complex types of nodes and edges, require specialized pre-training strategies. These can be broadly categorized into generative and contrastive methods. Generative learning in heterogeneous graphs primarily focuses on reconstructing masked segments of the input graph, either in terms of the underlying graph structures or specific node attributes (Hu et al., 2020b; Fang et al., 2022b; Tian et al., 2023). On the other hand, contrastive learning on heterogeneous graphs aims to refine node representations by magnifying the mutual information of positive pairs while diminishing that of negative pairs. Specifically, representations generated from the same data instance form a positive pair, while those from different instances constitute a negative pair. Some methods emphasizes contrasting node-level representations (Jiang et al., 2021a; Yang et al., 2022; Wang et al., 2021a; Jiang et al., 2021b), while another direction contrasts node-level representations with graph-level representations (Park et al., 2020; Jing et al., 2021; Ren et al., 2019). In general, the efficacy of contrastive methods surpasses that of generative ones (Tian et al., 2023), making them the default pre-training strategies adopted in this paper.
Prompt-based learning on graphs. The recent trend in Natural Language Processing (NLP) has seen a shift from traditional fine-tuning of pre-trained language models (LMs) to a new paradigm: “pre-train, prompt(Liu et al., 2023b). Instead of fine-tuning LMs through task-specific objective functions, this paradigm reformulates downstream tasks to resemble pre-training tasks by incorporating textual prompts to input texts. This not only bridges the gap between pre-training and downstream tasks but also instigates further research integrating prompting with pre-trained graph neural networks (Sun et al., 2023). For example, GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) introduce prompt templates to align the pretext task of link prediction with downstream classification. GPF (Fang et al., 2022a) and VNT-GPPE (Tan et al., 2023) employ learnable perturbations to the input graph, modulating pre-trained node representations for downstream tasks. However, all these techniques cater exclusively to homogeneous graphs, overlooking the distinct complexities inherent to the heterogeneity in real-world systems.
@@ 3. Preliminaries
Figure 1(a) illustrates an example heterogeneous graph with three types of nodes: author (A), paper (P), and subject (S), as well as two types of edges: “write” and “belong to”.
@@
Figure 1(b) presents the network schema for a heterogeneous graph. As per the network schema, we learn that a paper is written by an author and that a paper belongs to a subject.
@@
Figure 1(c) shows two metapaths for a heterogeneous graph: “PAP” represents that two papers are written by the same author, while “PSP” indicates that two papers share the same subject.
@@ 4.3. Heterogeneous Feature Prompt (C3)
Prompting techniques fundamentally revolve around the idea of augmenting the input data to better align with the pretext objectives. This makes the design of a graph-level transformation an important factor for the efficacy of prompting. To illustrate, let’s consider a homogeneous graph ${\mathcal{G}}$ with its adjacency matrix ${\bm{A}}$ and node feature matrix ${\bm{X}}$. We introduce $t_{\xi}$, a graph-level transformation function parameterized by $\xi$, such as changing node features, adding or removing edges, etc. Prior research (Fang et al., 2022a; Sun et al., 2023) has proved that for any transformation function $t_{\xi}$, there always exists a corresponding feature prompt ${\bm{p}}^{*}$ that satisfies the following property:
@@ 4.5. Prompt-Based Learning and Inference
where $\lambda$ is a regularization hyperparameter. The orthogonal regularization (Brock et al., 2016) loss ${\mathcal{L}}_{orth}$ is defined to ensure the label tokens in the virtual class prompt remain orthogonal during prompt tuning, fostering diversified representations of different classes:

### v4

@@ Abstract.
Heterogeneous graph neural networks; pre-training; prompt tuning
journalyear: 2024copyright: rightsretainedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645685isbn: 979-8-4007-0171-9/24/05ccs: Computing methodologies Neural networks
@@ 1. Introduction
One potential solution that could partially address these challenges is to adapt the “pre-train, prompt” paradigm from natural language processing (NLP) to the graph domain. In NLP, prompt-based learning has effectively generalized pre-trained language models across diverse tasks. For example, a sentiment classification task like “The Web Conference will take place in the scenic city of Singapore in 2024” can be reframed by appending a specific textual prompt “I feel so [MASK]” to the end. A language model pre-trained on next word prediction is highly likely to predict “[MASK]” as “excited” instead of “frustrated”, without necessitating extensive fine-tuning. With this methodology, certain downstream tasks can be seamlessly aligned with the pre-training objectives. While few prior works (Sun et al., 2023, 2022; Fang et al., 2023; Liu et al., 2023a; Tan et al., 2023) have delved into crafting various prompting templates for graphs, their emphasis remains strictly on homogeneous graphs. This narrow focus underscores the last challenge inherent to the heterogeneous graph structures typical of the Web:
@@
To comprehensively address all four aforementioned challenges, we propose HetGPT, a general post-training prompting framework tailored for heterogeneous graphs. Represented by the acronym Heterogeneous Graph Prompt Tuning, HetGPT serves as an auxiliary system for HGNNs that have undergone constrastive pre-training. At the core of HetGPT is a novel graph prompting function that reformulates the downstream node classification task to align closely with the pretext contrastive task. We begin with the virtual class prompt, which generalizes the intricate characteristics of each class in the embedding space. Then we introduce the heterogeneous feature prompt, which acts as a task-specific augmentation to the input graph. This prompt is injected into the feature space and the prompted node features are then passed through the pre-trained HGNN, with all parameters in a frozen state. Furthermore, a multi-view neighborhood aggregation mechanism, that encapsulates the complexities of the heterogeneous neighborhood structure, is applied to the target node, generating a node token for classification. Finally, Pairwise similarity comparisons are performed between the node token and the class tokens derived from the virtual class prompt via the contrastive learning objectives established during pre-training, which effectively simulates the process of deriving a classification decision. In summary, our main contributions include:
@@ 2. Related Work
Heterogeneous graph pre-training. Pre-training techniques have gained significant attention in heterogeneous graph machine learning, especially under the scenario with limited labeled nodes (Liu et al., 2022; Xie et al., 2022). Heterogeneous graphs, with their complex types of nodes and edges, require specialized pre-training strategies. These can be broadly categorized into generative and contrastive methods. Generative learning in heterogeneous graphs primarily focuses on reconstructing masked segments of the input graph, either in terms of the underlying graph structures or specific node attributes (Hu et al., 2020b; Fang et al., 2022; Tian et al., 2023). On the other hand, contrastive learning on heterogeneous graphs aims to refine node representations by magnifying the mutual information of positive pairs while diminishing that of negative pairs. Specifically, representations generated from the same data instance form a positive pair, while those from different instances constitute a negative pair. Some methods emphasizes contrasting node-level representations (Jiang et al., 2021a; Yang et al., 2022; Wang et al., 2021a; Jiang et al., 2021b), while another direction contrasts node-level representations with graph-level representations (Park et al., 2020; Jing et al., 2021; Ren et al., 2019). In general, the efficacy of contrastive methods surpasses that of generative ones (Tian et al., 2023), making them the default pre-training strategies adopted in this paper.
Prompt-based learning on graphs. The recent trend in Natural Language Processing (NLP) has seen a shift from traditional fine-tuning of pre-trained language models (LMs) to a new paradigm: “pre-train, prompt(Liu et al., 2023b). Instead of fine-tuning LMs through task-specific objective functions, this paradigm reformulates downstream tasks to resemble pre-training tasks by incorporating textual prompts to input texts. This not only bridges the gap between pre-training and downstream tasks but also instigates further research integrating prompting with pre-trained graph neural networks (Sun et al., 2023). For example, GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) introduce prompt templates to align the pretext task of link prediction with downstream classification. GPF (Fang et al., 2023) and VNT-GPPE (Tan et al., 2023) employ learnable perturbations to the input graph, modulating pre-trained node representations for downstream tasks. However, all these techniques cater exclusively to homogeneous graphs, overlooking the distinct complexities inherent to the heterogeneity in real-world systems.
@@ 4.3. Heterogeneous Feature Prompt (C3)
Prompting techniques fundamentally revolve around the idea of augmenting the input data to better align with the pretext objectives. This makes the design of a graph-level transformation an important factor for the efficacy of prompting. To illustrate, let’s consider a homogeneous graph ${\mathcal{G}}$ with its adjacency matrix ${\bm{A}}$ and node feature matrix ${\bm{X}}$. We introduce $t_{\xi}$, a graph-level transformation function parameterized by $\xi$, such as changing node features, adding or removing edges, etc. Prior research (Fang et al., 2023; Sun et al., 2023) has proved that for any transformation function $t_{\xi}$, there always exists a corresponding feature prompt ${\bm{p}}^{*}$ that satisfies the following property:
@@ 4.5. Prompt-based Learning and Inference
where $\lambda$ is a regularization hyperparameter. The orthogonal regularization (Brock et al., 2017) loss ${\mathcal{L}}_{orth}$ is defined to ensure the label tokens in the virtual class prompt remain orthogonal during prompt tuning, fostering diversified representations of different classes:
@@ 5. Experiments
  • (RQ1) Can HetGPT improve the performance of pre-trained heterogeneous graph neural networks on the semi-supervised node classification task?
  • (RQ2) How does HetGPT perform under different settings, i.e., ablated models and hyperparameters?
  • (RQ3) How does the prompt tuning efficiency of HetGPT compare to its fine-tuning counterpart?
  • (RQ4) How interpretable is the learned prompt in HetGPT?
5.1. Experiment Settings
5.1.1. Datasets
We evaluate our methods using three benchmark datasets: ACM (Zhao et al., 2020), DBLP (Fu et al., 2020), and IMDB (Fu et al., 2020). Detailed statistics and descriptions of these datasets can be found in Table 5. For the semi-supervised node classification task, we randomly select 1, 5, 20, 40, or 60 labeled nodes per class as our training set. Additionally, we set aside 1,000 nodes for validation and another 1,000 nodes for testing. Our evaluation metrics include Macro-F1 and Micro-F1.
5.1.2. Baseline models
We compare our approach against methods belonging to three different categories:
  • Supervised HGNNs: HAN (Wang et al., 2019), HGT (Hu et al., 2020a), MAGNN (Fu et al., 2020);
  • HGNNs with “pre-train, fine-tune”:
    • Generative: HGMAE (Tian et al., 2023);
    • Contrastive (our focus): DMGI (Park et al., 2020), HeCo (Wang et al., 2021a), HDMI (Jing et al., 2021);
  • GNNs with“pre-train, prompt: GPPT (Sun et al., 2022).
5.1.3. Implementation details
For the homogeneous method GPPT, we evaluate using all the metapaths and present the results with the best performance. Regarding the parameters of other baselines, we adhere to the configuration specified in their original papers.
In our HetGPT model, the heterogeneous feature prompt is initialized using Kaiming initialization (He et al., 2015). During the prompt tuning phase, we employ the Adam optimizer (Kingma and Ba, 2014) and search within a learning rate ranging from 1e-4 to 5e-3. We also tune the patience for early stopping from 20 to 100. The regularization hyperparameter $\lambda$ is set to 0.01. We experiment with the number of feature tokens $K$, searching values from { 1, 5, 10, 15, 20 }. Lastly, for our non-linear activation function $\sigma(\cdot)$, we use LeakyReLU.
5.2. Performance on Node Classification (RQ1)
Experiment results for semi-supervised node classification on three benchmark datasets are detailed in Table 5. Compared to the pre-trained DMGI, HeCo, and HDMI models, our post-training prompting framework, HetGPT, exhibits superior performance in 88 out of the 90 comparison pairs. Specifically, we observe a relative improvement of 3.00% in Macro-F1 and 2.62% in Micro-F1. The standard deviation of HetGPT aligns closely with that of the original models, indicating that the improvement achieved is both substantial and robust. It’s crucial to note that the three HGNNs with “pre-train, fine-tune” - DMGI, HeCo, and HDMI, are already among the state-of-the-art methods for semi-supervised node classification. By integrating them with HetGPT, we push the envelope even further, setting a new performance pinnacle. Furthermore, HetGPT’s edge becomes even more significant in scenarios where labeled nodes are extremely scarce, achieving an improvement of 6.60% in Macro-F1 and 6.88% in Micro-F1 under the 1-shot setting. Such marked improvements in few-shot performance strongly suggest HetGPT’s efficacy in mitigating the overfitting issue. The strategic design of our prompting function, especially the virtual class prompt, effectively captures the intricate characteristics of each class, which can potentially obviate the reliance on costly annotated data. Additionally, GPPT lags considerably on all datasets, which further underscores the value of HetGPT’s effort in tackling the unique challenges inherent to heterogeneous graphs.
5.3. Performance under Different Settings (RQ2)
5.3.1. Ablation study
To further demonstrate the effectiveness of each module in HetGPT, we conduct an ablation study to evaluate our full framework against the following three variants:
  • w/o VCP: the variant of HetGPT without the virtual class prompt from Section 4.2;
  • w/o HFP: the variant of HetGPT without the heterogeneous feature prompt from Section 4.3;
  • w/o MNA: the variant of HetGPT without the multi-view neighborhood aggregation from Section 4.4.
Experiment results on ACM and DBLP, shown in Figure 2, highlight the substantial contributions of each module to the overall effectiveness of HetGPT. Notably, the virtual class prompt emerges as the most pivotal component, indicated by the significant performance drop when it’s absent. This degradation mainly stems from the overfitting issue linked to the negative transfer problem, especially when labeled nodes are sparse. The virtual class prompt directly addresses this issue by generalizing the intricate characteristics of each class within the embedding space.
(a) ACM
(b) IMDB
Figure 2. Ablation study of HetGPT on ACM and IMDB.
5.3.2. Hyper-parameter sensitivity
We evaluate the sensitivity of HetGPT to its primary hyperparameter: the number of basis feature tokens $K$ in Equation (7). As depicted in Figure 3, even a really small value of $K$ (i.e., 5 for ACM, 20 for DBLP, and 5 for IMDB) can lead to satisfactory node classification performance. This suggests that the prompt tuning effectively optimizes performance without the need to introduce an extensive number of new parameters.
(a) ACM, DBLP
(b) IMDB
Figure 3. Performance of HetGPT with the different number of basis feature vectors on ACM, DBLP, and IMDB.
5.4. Prompt Tuning Efficiency Analysis (RQ3)
Our HetGPT, encompassing the virtual class prompt and the heterogeneous feature prompt, adds only a few new trainable parameters (i.e., comparable to a shallow MLP). Concurrently, the parameters of the pre-trained HGNNs and the contrastive head remain unchanged during the entire prompt tuning phase. Figure 4 illustrates that HetGPT converges notably faster than its traditional “pre-train, fine-tune” counterpart, both recalibrating the parameters of the pre-trained HGNNs and introducing a new prediction head. This further demonstrates the efficiency benefits of our proposed framework, allowing for effective training with minimal tuning iterations.
5.5. Interpretability Analysis (RQ4)
To gain a clear understanding of how the design of the virtual class prompt facilitates effective node classification without relying on the traditional classification paradigm, we employ a t-SNE plot to visualize the node representations and the learned virtual class prompt on ACM and DBLP, as shown in Figure 5. Within this visualization, nodes are depicted as colored circles, while the class tokens from the learned virtual class prompt are denoted by colored stars. Each color represents a unique class label. Notably, the embeddings of these class tokens are positioned in close vicinity to clusters of node embeddings sharing the same class label. This immediate spatial proximity between a node and its respective class token validates the efficacy of similarity measures inherited from the contrastive pretext for the downstream node classification task. This observation further reinforces the rationale behind our node classification approach using the virtual class prompt, i.e., a node is labeled as the class that its embedding is most closely aligned with.
(a) DBLP
(b) IMDB
Figure 4. Comparison of training losses over epochs between HetGPT and its fine-tuning counterpart on DBLP and IMDB.
(a) ACM
(b) DBLP
Figure 5. Visualization of the learned node tokens and class tokens in virtual class prompt on ACM and DBLP.
6. Conclusion
In this paper, we propose HetGPT, a general post-training prompting framework to improve the node classification performance of pre-trained heterogeneous graph neural networks. Recognizing the prevalent issue of misalignment between the objectives of pretext and downstream tasks, we craft a novel prompting function that integrates a virtual class prompt and a heterogeneous feature prompt. Furthermore, our framework incorporates a multi-view neighborhood aggregation mechanism to capture the complex neighborhood structure in heterogeneous graphs. Extensive experiments on three benchmark datasets demonstrate the effectiveness of HetGPT. For future work, we are interested in exploring the potential of prompting methods in tackling the class-imbalance problem on graphs or broadening the applicability of our framework to diverse graph tasks, such as link prediction and graph classification.